# Synthesizing data products, mathematical models, and observational measurements for lake temperature forecasting

### v1

@@ 3.3 GP surrogates and scalability
Beyond missing off-diagonal covariance terms, this is an inefficient estimator compared to the ideal SK quantity in Eq. (10) because it does not borrow strength nearby in the input space. (This is what the term $-\tau^{2}k(\mathbf{X}_{\_}n,\mathcal{X})^{\top}(\tau^{2}k(\mathbf{X}_{\_}n)+\mathbf{S}_{\_}n)^{-1}\tau^{2}k(\mathbf{X}_{\_}n,\mathcal{X})$ in that expression is doing.) So in $\sigma^{2(m)}_{\_}n(\mathbf{x})$ we’re getting over-inflated, conservative UQ.
@@
To benchmark our Vecchia-based workaround for SK we developed the following exercise to compare our SK approximation to true SK with respect to out of sample prediction accuracy and UQ. We used a subset of our corpus of NOAA-GLM runs. The goal was to have a large enough training exercise to stress our Vecchia calculations, while not being too large as to preclude expedient ordinary-GP calculations behind SK. We divided up one year’s worth of NOAA-GLM hindcasts, including all depths and horizons, into five-day chunks. For each chunk we used a random $n_{\_}i=16$ sample of NOAA-replicates for training, setting aside the other 15 for out-of-sample testing. In this way, each training data set was manageably sized at $n=1600$ unique inputs. For each of the seventy-three spans of five days in 2022 we fit SK (910) and the Vecchia-approximated first- and second-moments, $\mu_{\_}n^{(m)}$ and $\sigma^{2(m)}_{\_}n$ for the mean and $\mu^{(v)}_{\_}n$ for full predictive uncertainty in Eq. (12), respectively. We saved out-of-sample RMSEs and coverages of 95% CIs and PIs. The calculation of CI coverage involved averages of the fifteen held-out replicates, at each unique input, whereas PIs used all of the hold-out replicates individually.
[…]
We are encouraged by these results. True SK is not tractable on larger problems. Our Vecchia approximation can be fit on millions of NOAA-GLM runs in mere seconds, and we are confident that it closely matches the performance of SK.
3.4 Modeling forecast bias
Let $\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathbf{x})$ and $\hat{\sigma}^{\mathrm{SK}}(\mathbf{x})^{2}$ denote the moments of our NOAA-GLM surrogate (12), as described in Section 3.3 but fit to the entire campaign of runs across all 31 ensemble members and all training days/horizons/depths under study: ($\mathbf{X}_{\_}{n_{\_}M},\mathbf{Y}_{\_}{n_{\_}M}$). A fit based on these 8 million-odd runs, spanning three years of simulations, could only be fit via an approximation such as with Vecchia. Providing $\mathbf{x}\in\mathcal{X}$ representing any time point(s) of interest (future, present or past), including depths and horizons of interest, yields a prediction (or forecast) capturing uncertainty in GLM dynamics driven by NOAA ensemble hindcasts. Recall from Section 2.2 that estimating a bias correction requires measuring discrepancy between computer model predictions and field data observations, forming a data set $D^{b}$ that can be used to train a GP. Our field data are a single set of observations $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ in time and depth, coded in $\mathbf{X}_{\_}{n_{\_}F}$. In Section 2.2 we had $\mathbf{X}_{\_}{n_{\_}F}=\mathbf{X}_{\_}{n_{\_}M}$, so $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ could easily be lined up with predictions $\boldsymbol{\mu}_{\_}{n_{\_}M}$ to measure discrepancies (6). However, now our set of GLM inputs (and outputs) is expanded [Section 3.1] to encompass horizon, which is not relevant to our field sensor measurements. Field data and surrogate are out of alignment.
[…]
We may now define a data set of discrepancies, following Eq. (6), $D^{b}_{\_}{n}=(\bar{\mathbf{X}}_{\_}n,\bar{\mathbf{Y}}^{F}_{\_}{n_{\_}F}-\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathbf{X}_{\_}n))$, and fit $\mathcal{GP}(D^{b}_{\_}n)$ to characterize a bias correction over time/depth/horizon. The size of $n$ again makes full GP fitting prohibitive, so we use a Vecchia approximation. However, we do not have the luxury of replication, so we must opt for an ordinary, homoskedastic GP fit. Let $\mu_{\_}{n}^{b}(\cdot)$ and $\sigma_{\_}n^{2b}(\cdot)$ denote the moments of this GP fit, overloading our earlier Section 2.2 notation. A GP surrogate with bias-correction (GPBC) may be formed by chaining the two sets of predictive moments together.
$\mu_{\_}n(\mathcal{X})$ $\equiv\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathcal{X})+\mu_{\_}n^{b}(\mathcal{X})$ (13)
$\sigma_{\_}n^{2}(\mathcal{X})$ $\equiv\hat{\sigma}_{\_}n^{\mathrm{SK}}(\mathcal{X})^{2}+\sigma_{\_}n^{2b}(\mathcal{X})$
@@ 3.4 Modeling forecast bias
As in Section 2.2, $\sigma_{\_}n^{2}(\mathcal{X})$ over-estimates the variance of the sum of two GP variances because it is missing a negative covariance term. In contrast to the discussion in Section 2.2, we cannot utilize full predictive covariance matrices $\Sigma(\mathcal{X})$ because the Vecchia implementation we use only provides the diagonal, which is what we are denoting as $\sigma^{2}(\mathcal{X})$. We may put in for $\mathcal{X}$ anything we wish; however, the most useful settings indicate hindcasts ($\mathcal{X}$ depicting a historical set of forecasting days, horizons, and depths) and forecasts ($\mathcal{X}$ denoting the current day, a horizon of the next thirty days, and all depths of interest).
@@
Figure 8: NOAA GLM forecasts (gray lines) starting on Feb. 19 for years 2021, 2022, and 2023 at a depth of 1m. Sensor measurements (purple points) are overlaid. Row A: In-sample predictions and 90% PIs from the GLM surrogate (GLMGP, orange). Means and PIs for the bias GP (biasGP) in teal; raw bias in magenta. Row B: In-sample forecasts (predictive means) and 90% PIs for the bias-corrected GLM surrogate (GPBC, orange). Means and PIs for the bias GP in teal; and bias in magenta.
@@
To illustrate, Figure 8 shows examples of 1m-depth “slices” of hindcasts via quantities defined above starting on Feb. 19, 2021 (left), 2022 (middle) and 2023 (right). Consider the top row (denoted as A) first. Each gray line is a NOAA-GLM hindcast over a 30-day horizon. Orange lines summarize the computer model surrogate ($\hat{\mu}_{\_}n^{\mathrm{SK}}$ and $\hat{\sigma}_{\_}n^{\mathrm{SK}}$ via 90% error bars), denoted as GPGLM. [^4] We use the full PI here, via $\mu^{(v)}$, instead of $\sigma^{2(m)}$ providing a CI, which is used in the GPBC, as the former is easier to see. These are also what you would use if you were predicting via the surrogate alone, without bias correction. An example of the latter is shown in Figure 6. Purple dots show the $\bar{\mathbf{Y}}^{F}_{\_}{n_{\_}F}$ corresponding to that date/depth and each horizon. Discrepancy data $D^{b}$, derived from the $\hat{\mu}_{\_}n^{\mathrm{SK}}$ and $\bar{\mathbf{Y}}^{F}_{\_}{n_{\_}F}$, are shown at the bottom of each panel. Observe that in all three panels/years, these fits exhibit a cold bias because the bias correction is positive. The strength and statistical significance of this conclusion varies considerably across the three years. In 2021 correction from zero might not be statistically significant. But in 2022 and 2023 it is substantial (upwards of 3 degrees), highly non-linear, and error-bars do not cover zero over most of the 30-day horizon.
@@
The bottom row of the figure (denoted as B) shows how Eq. (3.4) is used to combine the quantities in the top row in order to provide the bias-corrected surrogate (GPBC). Observe that in all three years this provides a very good approximation to sensor measurements. PI coverage is excellent. Note that while PIs for the surrogate in the top/A panels exhibit a trumpet-like shape over increasing horizon, those in the bottom/B panels no longer have this pronounced shape. This is because our bias correction is homoskedastic.
@@
Figure 9: NOAA GLM forecasts (gray lines) starting on Feb. 19 for year 2023 at a depth of 1m; sensor measurements are shown in purple. Panel A shows the in-sample prediction from the GLM surrogate as depicted in 8 (right panel). Panel B shows the out-of-sample forecast from the bias-corrected GLM surrogate.
@@
Figure 9 shows a similar pair of views, but now in a forecasting context corresponding to the right-most column of Figure 8 (2023). Specifically, we mean that the gray lines, from the GLM were used to build the surrogate (along with simulations for all other days/depths in the past), but the field data (purple dots) are not observed. Thus, bias correction is based only on patterns learned on historical data. We chose this particular case to illustrate because it represents a challenging one. Although the overall predictive quality in Figure 9 is lower than the right column of Figure 8, because the former is in-sample and the later is out-of-sample, it is clear the the GPBC performs well. Visually, the bias correction is effective (panel B) compared to the surrogate on it’s own (panel A). This can also be quantified as follows, offered here as a precursor to a more exhaustive validation described next in Section 5: GPBC achieves an RMSE of 1.24C compared to 1.86C GPGLM on it’s own. Note that while PI coverage is lower for GPBC for this particular example, the width of the PIs of GPBC (2.23) is less than half of the PI width of the surrogate on its own (5.14).
@@ 4 Real-time forecasting
We now have all of the ingredients in place to build a real-time forecasting system combining simulated and observed temperatures. By “real-time,” we mean that forecasts can be updated daily, in a matter of minutes, when new field data (sensor measurements) and data products (NOAA forecasts) arrive. We have ensured that the process of fitting surrogates and bias correction is swift, so the primary delay comes from awaiting updated data. Here we elaborate on the configuration of our real-time system, with an emphasis on the principal variation based on Vecchia/SK, as described in Section 3, while also accommodating more straightforward alternatives earlier. We then describe a variation that we developed in order to allow forecasts to react more dynamically to the most recent observational and simulation data, and to data from the most similar previous years.
@@ 4.1 Iterative forecasting
Our forecasting framework consists of: an initial training stage, in which many historical NOAA-GLM simulations are used to fit a GP surrogate $\mathcal{GP}(D_{\_}{n_{\_}{M}})$ and bias correction $\mathcal{GP}(D^{b}_{\_}n)$; and an iterative stage, in which new data is downloaded, existing datasets are updated, and a new forecast is generated from the (updated) surrogates. Here we are using $\mathcal{GP}$ generically, since many variations are possible. The initial training phase begins with obtaining NOAA-GLM simulations (hindcasts) for all reference dates in the historical training period. A surrogate fit to these data provides a statistical abstraction of NOAA-GLM simulations over the historical period of time. Observed residuals relating sensor data to surrogate predictions may be used to estimate a bias correction.
The iterative phase (Fig. 10), which happens daily, is more interesting. This begins with downloading NOAA ensemble forecasts over the next thirty days. Next, these weather forecasts are fed as driver data to the GLM, augmenting the simulation campaign from the training phase, and updating the surrogate and bias correction accordingly. The bias-corrected surrogate can then be used to make lake temperature forecasts over that thirty-day period. The process wraps up, before repeating the next day, by observing actual temperatures on the current day from sensors at each depth. Our implementation is faithful to that description, but is more accurately described by an operation that is performed all at once in a matter of minutes at midnight (00UTC), separating “yesterday” from “tomorrow”. This is when NOAA forecasts for “tomorrow” (and the 29 subsequent days), and sensor measurements from “yesterday” become available. Yesterday’s sensor measurements can be used to validate previous forecasts, and the metrics we prefer are discussed next in Section 5.1. NOAA forecasts can be fed into GLM to produce 1–30 day-ahead lake temperature forecasts over each horizon and depth of interest. Those NOAA-GLM forecasts are then incorporated into the corpus of existing simulation training data, along with the sensor measurements. Model fitting for surrogates and bias correction can then again be carried out, for example as described in Sections 3.33.4. Although there is potential for computational economy in priming aspects of new fits from previous ones, we do not generally bother because the Vecchia-based methods are fast. Updated surrogates may then be used to generate a 1–30 day-ahead bias-corrected forecast. The process repeats again at the next midnight as part of the iterative, near-term forecasting cycle (Dietze et al., 2018).
@@ 4.2 Incorporating annual variability into forecasts
The inputs to our surrogate modeling and bias correction framework are Julian day $t\in\{1,\dots,365\}$, depth $d\in\{0,\dots,9\}$, and horizon $h\in\{1,\dots,30\}$. A consequence of this choice, especially with regard to the time input $t$, is that our forecasts will be identical for any day $t$ in any year. We have a “day of year” model, meaning that a forecast for today would be the same as a hindcast starting on that same day in any year represented in the training data or beyond. (Those hindcasts would differ, of course, from what we would have forecast at earlier times, as the training data of NOAA-GLM simulations and observations is augmented every time a forecast is made). As we gather more data, such “day of year” forecasts will have more conservative UQ because they will have been trained on a wider diversity of observations across year for that day, and days nearby in time. In other words, they will better capture climatological variability. But they will, as more time passes and data is gathered, be less reactive to current weather patterns and other environmental variables.
Figure 11: In-sample forecasts for March 1, 2023 from forecasting models without $\phi$ (“without $\phi$”) and those with $\phi$ (“with” $\phi$) at 1m depth; 90% prediction intervals are denoted by dashed lines. Sensor measurements are shown in purple. GLM forecasts for March 1, 2023 are shown in dark gray; previous years are shown in light gray.
To illustrate, consider the left column of Figure 11, which shows examples of predictions over a 30-day horizon for March 3 2023 at depth 1m from the surrogate only (top panel) and after bias correction (bottom). As with similar figures shown previously, forecasts are in orange (solid mean and dashed intervals), with sensor measurements as purple dots. In contrast to previous visuals, we now show all three sets of NOAA-GLM simulations available up until this time point: today’s (March 3 2023) in black and the previous two years (March 3 in 2022 and 2021) in gray. Observe that the orange forecasts do a poor job of tracking with today’s black simulations. In the top row, forecasts from the surrogate track the corpus of simulations, whereas in the bottom row their spread is narrowed somewhat by the bias correction, but in an overly conservative fashion.
Toward obtaining more dynamic forecasts, we introduce a fourth input variable, which we call $\phi$, whose behavior is illustrated in the right panels in the figure and is described as follows. With $\phi$ we aim to capture the unique “state” for day $t$ and year $\lambda$ [first introduced in Section 3.1], via the following calculation
@@
capturing the average sensor-measured temperatures near the surface (depths zero and one meters) over the previous five days. We then augment design matrices $\bar{\mathbf{X}}_{\_}n$ and $\bar{\mathbf{X}}_{\_}{n_{\_}F}^{F}$ to include $\phi(t,\lambda)$ matching up with $t$ and $\lambda$ already recorded in its rows. (Recall that the $\lambda$ column is ignored in our GP modeling, but now with the addition of $\phi$ its presence is felt through the state that it captures (14). Surrogate modeling and bias correction, as described in Sections 3.23.4, are not affected by this change excepting the higher input dimension, and corresponding additional lengthscale(s). In particular, responses $\bar{\mathbf{Y}}_{\_}n^{M}$ and $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ remain as before.
To see the effect of adding $\phi$ as a predictor consider the right column of Figure 11. Observe how, in contrast with the left column without $\phi$, the GLM surrogate without a bias-correction (top/orange) provides a good fit to NOAA-GLM forecasts (black lines). When bias-correction is added (bottom/orange), prediction intervals narrow considerably and appropriately compared to the sensor measurements (purple dots). We generally find better results with the addition of $\phi$, as we shall demonstrate momentarily in Section 5. Consequently, we shall going forward take the $\phi$ version to be the main/default modeling variation unless we mention that modeling “without $\phi$” was entertained as a contrast. In fact, all previous visuals in the paper used the version with $\phi$, because they looked better. The diagram in Figure 10 lists $\phi$ as one of the inputs. Of course, Eq. (14) is only one of possibly many ways that a year-differentiating state could be calculated. It could even be treated as a latent variable, to be inferred along with other unknowns, which we discuss as a possible extension in Section 6.

### v2

@@ 3.3 GP surrogates and scalability
Beyond missing off-diagonal covariance terms, this is an inefficient estimator compared to the ideal SK quantity in Eq. (10) because it does not borrow strength nearby in the input space. (This is what the term $-\tau^{2}k(\mathbf{X}_{\_}n,\mathcal{X})^{\top}(\tau^{2}k(\mathbf{X}_{\_}n)+\mathbf{S}_{\_}n)^{-1}\tau^{2}k(\mathbf{X}_{\_}n,\mathcal{X})$ in that expression is doing.) So in $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ we are getting over-inflated, conservative UQ because the Veccia software we are using does not provide a full covariance.
Appendix A.2A.3 provides two examples showing that this approach achieves the desired results in a benchmark/toy example and on GLM runs, respectively. These are intended to augment other visuals in the paper, which are all derived from Vechhia approximations.
3.4 Modeling forecast bias
Now, let $\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathbf{x})$ and $\hat{\sigma}^{\mathrm{SK}}(\mathbf{x})^{2}$ denote the moments of that NOAA-GLM surrogate (Eq. 11), combining Sections 3.23.3, but fit to the entire campaign of runs across all 31 ensemble members and all training days/horizons/depths under study: ($\mathbf{X}_{\_}{n_{\_}M},\mathbf{Y}_{\_}{n_{\_}M}$), and $n_{\_}M\approx$ 8 million-odd runs. Providing $\mathbf{x}\in\mathcal{X}$ representing any time point(s) of interest (future, present or past), including depths and horizons of interest, yields a prediction (or forecast) capturing uncertainty in GLM dynamics driven by NOAA ensemble hindcasts. Recall from Section 2.2 that estimating a bias correction requires measuring discrepancy between computer model predictions and field data observations, forming a data set $D^{b}$ that can be used to train a GP. Our field data are a single set of observations $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ in time and depth, coded in $\mathbf{X}_{\_}{n_{\_}F}$. In Section 2.2 we had $\mathbf{X}_{\_}{n_{\_}F}=\mathbf{X}_{\_}{n_{\_}M}$, so $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ could easily be lined up with predictions $\boldsymbol{\mu}_{\_}{n_{\_}M}$ to measure discrepancies (6). However, now our set of GLM inputs (and outputs) is expanded [Section 3.1] to encompass horizon, which is not relevant to our field sensor measurements. Field data and surrogate are out of alignment.
@@ 3.4 Modeling forecast bias
$\mu_{\_}n(\mathcal{X})$ $\equiv\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathcal{X})+\mu_{\_}n^{b}(\mathcal{X})$ and $\sigma_{\_}n^{2}(\mathcal{X})$ $\equiv\hat{\sigma}_{\_}n^{\mathrm{SK}}(\mathcal{X})^{2}+\sigma_{\_}n^{2b}(\mathcal{X}).$ (12)
@@
Recall that $\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathcal{X})$ is the predictive mean of the first moments’ fit $\mathcal{GP}(D_{\_}n^{(m)})$. As in Section 2.2, $\sigma_{\_}n^{2}(\mathcal{X})$ over-estimates the variance of the sum of two GP variances because it is missing a negative covariance term. In contrast to the discussion in Section 2.2, we cannot utilize full predictive covariance matrices $\Sigma(\mathcal{X})$ because the Vecchia implementation we use only provides the diagonal, which is what we are denoting as $\sigma_{\_}n^{2}(\mathcal{X})$. We may put in for $\mathcal{X}$ anything we wish; however, the most useful settings indicate hindcasts ($\mathcal{X}$ depicting a historical set of forecasting days, horizons, and depths) and forecasts ($\mathcal{X}$ denoting the current day, a horizon of the next thirty days, and all depths of interest).
To summarize, the NOAA-GLM surrogate arises from fitting GPs to first and second moments, $\bar{\mathbf{Y}}_{\_}n$ and $\bar{\mathbf{S}}_{\_}n$, respectively, via methods described in Sections 3.2 and 3.3. The predictive mean, $\hat{\mu}_{\_}n^{\mathrm{SK}}(\mathcal{X})$ is subtracted from observations $\bar{\mathbf{Y}}^{F}_{\_}n$ to form a dataset of discrepancies $D^{b}_{\_}n$, to which a homoscedastic GP is fitted. Finally, the GP surrogate with bias-correction (GPBC) is obtained by adding predictive moments from both the surrogate and bias correction.
@@
Figure 6: NOAA GLM forecasts (thin gray lines) starting on Feb. 19 for years 2021, 2022, and 2023 at a depth of 1m. Sensor measurements (dark points) are overlaid. Top row: In-sample predictions and 90% PIs from the GLM surrogate (GPGLM, thick lines). Middle row: Means and 90% PIs for the bias GP (biasGP) with bias values overlaid as points. Bottom row: In-sample forecasts (predictive means) and 90% PIs for the bias-corrected GLM surrogate (GPBC, thick lines) with observations overlaid as dark points.
@@
To illustrate, Figure 6 shows examples of 1m-depth “slices” of hindcasts via quantities defined above starting on Feb. 19, 2021 (left), 2022 (middle) and 2023 (right). Consider the top row first. Each thin gray line is a NOAA-GLM hindcast over a 30-day horizon. Thicker, orange lines summarize the computer model surrogate ($\hat{\mu}_{\_}n^{\mathrm{SK}}$ and $\hat{\sigma}_{\_}n^{\mathrm{SK}}$ via 90% error bars), denoted as GPGLM. Darker dots show the $\bar{\mathbf{Y}}^{F}_{\_}{n_{\_}F}$ corresponding to that date/depth and each horizon. Discrepancy data $D^{b}$, derived from the $\hat{\mu}_{\_}n^{\mathrm{SK}}$ and $\bar{\mathbf{Y}}^{F}_{\_}{n_{\_}F}$, are shown at the bottom of each panel. Observe that in all three panels/years, these fits exhibit a cold bias because the bias correction is positive. The strength and statistical significance of this conclusion varies considerably across the three years. In 2021 correction from zero might not be statistically significant. But in 2022 and 2023 it is substantial (upwards of 3 degrees), highly non-linear, and error-bars do not cover zero over most of the 30-day horizon.
@@
The bottom row of the figure shows how Eq. (12) is used to combine the quantities in the top and middle row, showing the bias fit, to obtain a bias-corrected surrogate (GPBC). Observe that in all three years this is a good approximation to sensor measurements. Note that while PIs for the surrogate in the top panels exhibit a trumpet-like shape over increasing horizon, those in the bottom panels no longer have this pronounced shape. GPBC is more confident than GPGLM because the latter uses $\mathbb{V}\mathrm{ar}\{Y(\mathbf{\mathbf{x}})\}$ but the former leverages $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ from Eq. (11). The total variance of GPBC is obtained by adding $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}$ to the predictive variance from the bias GP $(\sigma_{\_}n^{2b}(\mathcal{X}))$, shown in Eq. (12). Because $\hat{\sigma}^{\mathrm{SK}}_{\_}n(\mathbf{x})^{2}\ll\sigma_{\_}n^{2b}(\mathcal{X})$ from the (homoskedastic) bias GP, the overall predictive variance of GPBC appears homoscedastic.
@@
Figure 7: NOAA GLM forecasts (gray lines) starting on Feb. 19 for year 2023 at a depth of 1m; sensor measurements are shown as darker dots. Panel A shows the in-sample prediction from the GLM surrogate as depicted in 6 (right panel). Panel B shows the out-of-sample forecast from the bias-corrected GLM surrogate.
@@
Figure 7 shows a similar pair of views, but now in a forecasting context corresponding to the right-most column of Figure 6 (2023). Specifically, we mean that the gray lines, from the GLM were used to build the surrogate (along with simulations for all other days/depths in the past), but the field data (dark dots) are not observed. Thus, bias correction is based only on patterns learned on historical data. Although the overall predictive quality in Figure 7 is lower than the right column of Figure 6, because the former is in-sample and the later is out-of-sample, it is clear the GPBC performs well. Visually, the bias correction is effective (panel B) compared to the surrogate on its own (panel A). This can also be quantified as follows: GPBC achieves an RMSE of 1.24C compared to 1.86C GPGLM on its own. Note that while PI coverage is lower for GPBC for this particular example, the width of the PIs of GPBC (2.23) is less than half of the PI width of the surrogate on its own (5.14).
@@ 4 Real-time forecasting
We now have all of the ingredients in place to build a real-time forecasting system combining simulated and observed temperatures. Here we elaborate on the configuration of our system, with an emphasis on the principal variation based on Vecchia/SK, as described in Section 3, while also accommodating more straightforward alternatives earlier. We then describe a variation that allows forecasts to react more dynamically to recent observational and simulation data, and to data from the most similar previous years.
@@ 4.1 Iterative forecasting
Our forecasting framework consists of: an initial training stage, in which many historical NOAA-GLM simulations are used to fit a GP surrogate $\mathcal{GP}(D_{\_}{n_{\_}{M}})$ and bias correction $\mathcal{GP}(D^{b}_{\_}n)$; and an iterative stage, in which new data are downloaded and a new forecast is generated from updated surrogates. Here we are using $\mathcal{GP}$ generically, since many variations are possible. The initial training phase begins with obtaining NOAA-GLM blue hindcasts for all reference dates in the historical training period. A surrogate fit to these data provides a statistical abstraction of NOAA-GLM simulations over the historical period of time. Observed residuals from sensor data are be used to estimate a bias correction. The iterative phase, which happens daily, is more interesting. This begins with downloading NOAA ensemble forecasts over the next thirty days. Next, these are fed as driver data to the GLM, augmenting the simulation campaign from the training phase, and updating the surrogate and bias correction accordingly. The bias-corrected surrogate can then be used to make lake temperature forecasts over that thirty-day period. The process repeats the next day after observing actual temperatures on the current day from sensors at each depth. Additional detail, along with a diagrammatic depiction [Figure 14] is provided in Appendix B.
@@ 4.2 Incorporating annual variability into forecasts
The inputs to our surrogate modeling and bias correction framework are Julian day $t\in\{1,\dots,365\}$, depth $d\in\{0,\dots,9\}$, and horizon $h\in\{1,\dots,30\}$. A consequence of this choice, especially with regard to the time input $t$, is that our forecasts will be identical for any day $t$ in any year. We have a “day of year” model. As we gather more data, forecasts will have more conservative UQ because they will have been trained on a wider diversity of observations across year for that day, and days nearby in time. In other words, they will better capture climatological variability. But they will also be less reactive to current weather patterns and other environmental variables.
Figure 8: In-sample forecasts for March 1, 2023 models without $\phi$ (“without $\phi$”; left panels, dotted lines) and those with $\phi$ (“with” $\phi$); right panels, solid lines) at 1m depth; 90% PIS are denoted by dashed lines, Sensor measurements as purple dots. GLM forecasts for March 1, 2023 are in dark gray; previous years in light gray.
To illustrate, consider the left column of Figure 8, which shows examples of predictions over a 30-day horizon for March 1, 2023 at depth 1m from the surrogate only (top panel) and after bias correction (bottom). In contrast to previous visuals, we now show all three sets of NOAA-GLM simulations available up until this time point: today’s (March 1 2023) in black and the previous two years (March 1 in 2022 and 2021) in gray. Observe that the forecasts do a poor job of tracking today’s black simulations. In the top row, forecasts from the surrogate track the corpus of simulations, whereas in the bottom row their spread is narrowed by the bias correction, but in an overly conservative fashion.
Toward obtaining more dynamic forecasts, we introduce a fourth input variable, $\phi$, whose behavior is illustrated in the right panels in the figure. With $\phi$ we aim to capture the unique “state” for day $t$ and year $\lambda$ [first introduced in Section 3.1], via the following calculation
@@
i.e., average sensor-measured temperatures near the surface (depths $d$ = 0 and 1 meters) over the previous five days. Temperatures near the surface are more variable than those at deeper layers, and are therefore a good indication of weather conditions. We augment design matrices $\bar{\mathbf{X}}_{\_}n$ and $\bar{\mathbf{X}}_{\_}{n_{\_}F}^{F}$ to include $\phi(t,\lambda)$ matching up with $t$ and $\lambda$. (Recall that the $\lambda$ column is ignored in our GP modeling, but now with the addition of $\phi$ its presence is felt through the state that it captures (Eq. 13)). Surrogate modeling and bias correction, as described in Sections 3.23.4, are not affected by this change excepting higher input dimension and additional lengthscale(s). In particular, responses $\bar{\mathbf{Y}}_{\_}n^{M}$ and $\mathbf{Y}^{F}_{\_}{n_{\_}F}$ remain as before.
The right column of Figure 8 explores the effect of adding $\phi$ as a predictor. Observe how, in contrast with the left column without $\phi$, the GLM surrogate without a bias-correction (top/orange) provides a good fit to NOAA-GLM forecasts (black lines). When bias-correction is added (bottom/orange), prediction intervals narrow considerably compared to the sensor measurements (purple dots). We generally find better results with the addition of $\phi$, as demonstrated in Section 5. Consequently, we shall going forward take the $\phi$ version to be the main/default modeling variation unless we mention that modeling “without $\phi$” was entertained as a contrast. In fact, all previous visuals in the paper used the version with $\phi$, because they looked better. Eq. (13) represents only one of many ways that a year-differentiating state could be calculated. It could even be treated as a latent variable [see Appendix E].